# Maria Bazotte, Margarida Carvalho, and Thibaut Vidal

### v3

@@ 2.2 Literature Review
For discrete random variables, the random elements are usually assumed to be independent. For the discrete selection of distributions, distinct applications of two-stage programs have been studied, such as network design protection (Viswanath et al. 2004, Flach 2010, Peeta et al. 2010, Prestwich et al. 2014), facility protection (Medal et al. 2016), facility location-protection (Li et al. 2021), fire risk management (Bhuiyan et al. 2019), supplier risk management (Zhou et al. 2022), facility protection and network design (Bhuiyan et al. 2020), and airline flight network expansion problems (Şafak et al. 2022). All these works used decision-dependent probabilities as in Equation (4), except for Şafak et al. (2022), who chose the probabilities of the entire random vector. These applications usually lead to large-scale instances with numerous scenarios.
@@
Prestwich et al. (2014) and Bhuiyan et al. (2019) respectively obtained exact MILP reformulations of the DEP in the extensive form of the original program by using a linear scaling procedure and the probability chain technique described by O’Hanley et al. (2013). To solve larger-scale problems, Prestwich et al. (2014) proposed a bundle method, and Medal et al. (2016), Bhuiyan et al. (2020), and Zhou et al. (2022) implemented the L-shaped decomposition for solving reformulations resulting from the probability chain technique. Usually, $T_{k}$ is a zero matrix in these two-stage programs (i.e., the second stage can be solved a priori). The only exception is Zhou et al. (2022), who used an exact McCormick linearization (McCormick 1976) to obtain an MILP. These reformulations apply specifically to the case of discrete selection of distributions and rely on the binary variables in Equation (4). However, they require numerous variables and big-M constraints. Thus, the practical use of these reformulations is limited to problems with a sufficiently small number of scenarios. Flach (2010) proposed a different approximate MILP reformulation for the DEP by using the binary variables in Equation (4) to reformulate $p_{k}(x)$ as a set of exponential and linear constraints. The authors ensured a predefined precision by approximating the exponential constraints by piecewise linear cuts included in a branch-and-cut scheme.
@@
Various decision-dependent probability mass functions have also been studied, such as the binomial distribution (Karaesmen and Van Ryzin 2004, Green et al. 2013), linear interpolation (Du and Peeta 2014), a linear scaling and convex combination of a set of distributions (Hellemo et al. 2018), and the exponential function (Bhuiyan 2021). In most studies, the endogenous uncertainty in general cases of discrete random variables resulted in DEPs, which are mixed-integer nonlinear nonconvex programs (MINLPs) for which proving global optimality is harder than for MILPs or mixed-integer convex programs (MICPs). Even if $p_{k}(x)$ are convex or concave (see, e.g., Hellemo et al. 2018) or $\hat{p}_{\hat{r}(u,k)}^{u}(x)$ are convex or concave (see, e.g., Du and Peeta 2014, Bhuiyan 2021, Holzmann and Smith 2021, Homem-de Mello et al. 2022), the term $p_{k}(x)q_{k}^{\top}y_{k}$ could lead to an MINLP. Within this context, Krasko and Rebennack (2017) studied a two-stage preparedness hazard problem for debris-flow-hazard mitigation and solved the DEP with the BARON solver. Hellemo et al. (2018) solved the DEP of an energy investment problem using a generalized Benders decomposition (GBD) and the BARON solver. However, these approaches are limited to solving instances with few scenarios because the DEP in extensive form explicitly enumerates all scenarios.
@@
Viswanath et al. (2004), Peeta et al. (2010), and Du and Peeta (2014) used the SAA within iterative heuristic procedures to evaluate current first-stage solutions for two-stage network protection problems. Unfortunately, these procedures do not provide performance guarantees. The SAA was also used to estimate the performance of current solutions within stochastic gradient methods (Karaesmen and Van Ryzin 2004, Homem-de Mello et al. 2022) for overbooking and appointment-scheduling problems, respectively. Homem-de Mello et al. (2022) proved the local convergence of the algorithm, which is compatible with its nonconvexity. Finally, by considering a specific gradient estimator, Karaesmen and Van Ryzin (2004) proved the local convergence of the algorithm.
@@
To the best of our knowledge, only Ernst and Powell (1995), Ekin et al. (2017), Ekin (2018), and Hellemo et al. (2018) investigated continuous endogenous random variables. Ernst and Powell (1995) proposed a search procedure for solving a newsvendor problem with uncertain demand following a normal distribution with mean and variance selected by first-stage decisions. The augmented probability simulation (Bielza et al. 1999), in which simulated decisions are considered to be random variables, was used by Ekin et al. (2017) and Ekin (2018) for solving, respectively, a newsvendor problem with uncertain demand depending on the definition of the stock level and a production-planning problem with random yield depending on maintenance and production decisions. In both works, the dependency was modeled with a normal distribution with mean and variance defined according to the first-stage decisions. Finally, Hellemo et al. (2018) used the discretized function of the Kumaraswamy distribution and approximated the discretized function of the normal distribution. The discretization trades an increase in the number of scenarios for a confident discrete approximation. Approximations of the discretized function may also be needed when the cumulative distribution has no closed form or has a complex closed form. In some applications, the search procedure cannot analyze the entire decision space, implying the need for a heuristic procedure. Finally, the augmented probability simulation remains limited to problems such as the newsvendor problem where the second stage is represented by analytical expressions.
@@ 3 Methodology
Our methodology is based on using transformations of random variables, which are a classical topic in probability and statistics (see Hogg et al. 1977, Chap. 5), to model endogenous uncertainty. This section outlines Assumption 3, which is necessary for applying our methodology. Section 4 shows that this assumption is not restrictive by evidencing many cases in which it holds. {assumption} There exists (1) a random vector $\vartheta^{\top}(\omega)=\boldsymbol{(}\widetilde{q}(\omega)^{\top},\widetilde{h}(\omega)^{\top},\widetilde{T}(\omega),\widetilde{W}(\omega)\boldsymbol{)}$ with a known probability distribution that is independent of the first-stage decisions $x\in\mathcal{X}$ and (2) a transformation function $t(x,\boldsymbol{\vartheta})$, such that $\boldsymbol{\xi}_{x}=t\left(x,\boldsymbol{\vartheta}\right)$. Assumption 3 allows us to mimic the endogenous uncertainty by using a function $t(\cdot,\cdot)$ that combines the first-stage variables $x$ and the decision-independent random vector $\boldsymbol{\vartheta}$. For example, $t(\cdot,\cdot)$ can be the inverse transformation of $\boldsymbol{\xi}_{x}$ (Robert and Casella 1999, Chap. 2). Here, the random vector $\boldsymbol{\vartheta}$ does not necessarily have the same dimension as $\boldsymbol{\xi}$ or $\boldsymbol{\xi}_{x}$. Furthermore, the random elements of $\boldsymbol{\xi}_{x}$ may be dependent. For the sake of simplicity, we consider that $t(x,\boldsymbol{\vartheta})=(t_{1}\boldsymbol{(}x,\widetilde{q}(\omega)^{\top}\boldsymbol{)},t_{2}\boldsymbol{(}x,\widetilde{h}(\omega)\boldsymbol{)},t_{3}\boldsymbol{(}x,\widetilde{T}(\omega)\boldsymbol{)},t_{4}\boldsymbol{(}x,\widetilde{W}(\omega)\boldsymbol{)})$, where $t_{1}\boldsymbol{(}x,\widetilde{q}\left(\omega\right)^{\top}\boldsymbol{)}$, $t_{2}\boldsymbol{(}x,\widetilde{h}\left(\omega\right)\boldsymbol{)}$, $t_{3}\boldsymbol{(}x,\widetilde{T}\left(\omega\right)\boldsymbol{)}$, and $t_{4}\boldsymbol{(}x,\widetilde{W}\left(\omega\right)\boldsymbol{)}$ are transformation functions applied to the corresponding subvectors and matrices of $\boldsymbol{\vartheta}$. These functions return, respectively, the subvectors and matrices of $\boldsymbol{\xi}_{x}$, i.e., $q_{x}(\omega)^{\top}$, $h_{x}(\omega)$, $T_{x}(\omega)$, and $W_{x}(\omega)$. From Assumption 3, $t(.,.)$ and thus $t_{1}(.,.)$, $t_{2}(.,.)$, $t_{3}(.,.)$, and $t_{4}(.,.)$, must account for any dependency between the random elements of $\boldsymbol{\xi}_{x}$. Section 4 details different methods to define appropriate transformations respecting Assumption 3.
@@
In the field of stochastic optimization, Varaiya and Wets (1989) and Dupacová (2006) introduced the idea of transforming a decision-dependent probability distribution into an independent distribution in what is called “plug-in” or “change of measure” (see Pflug 2012 for a detailed explanation) and might employ random variable transformation. Despite these initial works, no study has directly modeled endogenous uncertainty as a function $t(.,.)$ nor investigated appropriate transformations for random variables following classical endogenous probability distributions. Specifically, the definition of $t(.,.)$ remains unexplored. Moreover, no study has focused on the nature and solution of problems resulting from this transformation nor, notably, addressed their potential nonconvexity. Finally, Holzmann and Smith (2021) used random variable transformation for a specific endogenous discrete random variable, for which Section 4.2.1 highlights the similarities and differences with the approach proposed herein. Therefore, the present work proposes a general framework using functions of random variables for addressing 2-SMIPD given Assumption 3. We also investigate practical transformations for endogenous random variables following classical probability distributions, as well as the convexity of the resulting transformed program.
Based on Assumption 3, we can formulate 2-SMIPD as
$\min$ $c^{\top}x+\mathbb{E}_{\boldsymbol{\vartheta}}\left[Q_{T}\boldsymbol{(}x,\vartheta(\omega)\boldsymbol{)}\right]$ (6a)
@@
[…]
$t_{3}\boldsymbol{(}x,\widetilde{T}\left(\omega\right)\boldsymbol{)}x+t_{4}\boldsymbol{(}x,\widetilde{W}\left(\omega\right)\boldsymbol{)}y(\omega)\geq t_{2}\boldsymbol{(}x,\widetilde{h}\left(\omega\right)\boldsymbol{)},y(\omega)\in\mathcal{Y}\},$
@@
Under Assumption 3, $g(x)=\mathbb{E}_{\boldsymbol{\xi}\mid x}\left[Q\boldsymbol{(}x,\xi(\omega)\boldsymbol{)}\right]=\mathbb{E}_{\boldsymbol{\vartheta}}\left[Q_{T}\boldsymbol{(}x,\vartheta(\omega)\boldsymbol{)}\right]$ and the 2-SMIPD (original Program (1)) can be reformulated into the transformed Program (6). In other words, Programs (1) and (6) have the same optimal solutions.
Thus, our methodology is established on finding an appropriate transformation function to obtain Program (6b). In that way, under Assumption 3, 2-SMIPD can be transformed into a two-stage stochastic program with exogenous uncertainty. Solution procedures for the transformed Program (6) based on its properties are well-studied (see, e.g., Birge and Louveaux 2011, Shapiro et al. 2021). Program (6b) has the terms $t_{1}\boldsymbol{(}x,\widetilde{q}\left(\omega\right)^{\top}\boldsymbol{)}y$, $t_{2}\boldsymbol{(}x,\widetilde{h}\left(\omega\right)\boldsymbol{)}$, $t_{3}\boldsymbol{(}x,\widetilde{T}\left(\omega\right)\boldsymbol{)}x$, and $t_{4}\boldsymbol{(}x,\widetilde{W}\left(\omega\right)\boldsymbol{)}y$, which might be nonlinear nonconvex. We then introduce the following assumption: {assumption} The continuous relaxation of the second-stage Program (6b) is convex. Program (6b) might need to be reformulated to convexify its continuous relaxation. Section 4 presents different cases for which Assumption 3 holds. In addition, the exogenous random vector $\boldsymbol{\vartheta}$ might have a large or infinite dimension (see, e.g., the inverse transformation in Section 4.2). Hence, next, we describe the application of the SAA to solve Program (6), which, although intuitive, shows that the transformation does not compromise desirable SAA properties. Nonetheless, other procedures can also be employed, e.g., stochastic gradient (Birge and Louveaux 2011).
@@ 3.1 Sample Average Approximation
We can solve Program (6) with the SAA (Shapiro et al. 2021, Chap. 5). We have $N$ independent and identically distributed samples $(\boldsymbol{\vartheta^{1}},\ldots,\boldsymbol{\vartheta^{N}})$ of $\boldsymbol{\vartheta}$ with realizations $(\vartheta^{1},\ldots,\vartheta^{N})$. The SAA program is
@@
where $\hat{g}_{N}(x)$ is the SAA estimator of $\mathbb{E}_{\boldsymbol{\vartheta}}\left[Q_{T}\boldsymbol{(}x,\vartheta(\omega)\boldsymbol{)}\right]$ defined as
@@
This estimator converges point-wise with probability one (w.p.1) to $\mathbb{E}_{\boldsymbol{\vartheta}}[Q_{T}\boldsymbol{(}x,\vartheta(\omega)\boldsymbol{)}]$ when $N\rightarrow+\infty$. From Lemma 3.1, we have $\mathbb{E}[\hat{g}_{N}(x)]=g(x)$, so $\hat{g}_{N}(x)$ is an unbiased estimator of $g(x)$. We denote $v^{*}$, $v_{t(x,\boldsymbol{\vartheta})}^{*}$ and $v_{N}$ as the optimal values of Program (1), Program (6), and Program (7), respectively. Under some mild conditions, such as when Program (6b) is convex, an MILP or when the first-stage feasible set $\mathcal{X}$ is finite, $v_{N}$ converges to $v_{t(x,\boldsymbol{\vartheta})}^{*}$ w.p.1 as $N\rightarrow+\infty$ (cf. Shapiro et al. 2021, Chap. 5, Ahmed and Shapiro 2002, and Kleywegt et al. 2002, respectively). In these cases and under Assumption 3, $v_{N}$ converges to $v^{*}$ w.p.1 as $N\rightarrow+\infty$, since $v^{*}=v_{t(x,\boldsymbol{\vartheta})}^{*}$ from Lemma 3.1. This result depends on the properties of $\boldsymbol{\vartheta}$ (i.e., on the convergence of $v_{N}$ to $v_{t(x,\boldsymbol{\vartheta)}}$), which may have larger dimension and support than $\boldsymbol{\xi}$ and may require a larger number of samples.
@@
The SAA finds a feasible first-stage solution $\overline{x}\in\mathcal{X}$. To obtain $\overline{x}$, we solve $M$ instances of the SAA Program (7), each with $N$ independent samples. We then evaluate the optimal solutions of every instance by using a larger sample size $N^{\prime}$ and define $\overline{x}$ as the optimal solution with the best evaluation. When $Q\boldsymbol{(}x,\xi(\omega)\boldsymbol{)}$ is finite and Program (1) has relatively complete recourse (i.e., the second stage associated with $x\in\mathcal{X}$ is feasible for almost every realization of $\boldsymbol{\xi}$), then $Q_{T}\boldsymbol{(}x,\vartheta(\omega)\boldsymbol{)}$ is also finite, and Program (6) also has relatively complete recourse, given that $\boldsymbol{\xi}_{x}=t(x,\boldsymbol{\vartheta})$ from Assumption 3. In this case, in addition to finding $\overline{x}$, we can compute a lower-bound estimator $\overline{v}_{M}^{N}$ of $v^{*}$, an upper-bound estimator $\hat{v}_{N^{\prime}}(\overline{x})$ of $v^{*}$, an over-estimator $\overline{gap}(\overline{x})$ of the optimality gap $c^{\top}\overline{x}+g(\overline{x})-v^{*}$, and unbiased estimators of the variances of $\overline{v}_{N}^{M}$, $\hat{g}_{N{{}^{\prime}}}(\overline{x})$, and $\overline{gap}(\overline{x})$, i.e., $\hat{\sigma}_{N,M}^{2}$, $\hat{\sigma}_{N{{}^{\prime}}}^{2}(\overline{x})$, and $\hat{\sigma}_{gap}^{2}(\overline{x})$, respectively. Appendix 8 details these estimators and the SAA’s procedure.
@@
By defining the $\boldsymbol{\xi}_{x}$ with the function $t(x,\boldsymbol{\vartheta})$, the estimators $\hat{g}_{N}(x)$ for all $x\in\mathcal{X}$ are based on the same samples of $\boldsymbol{\vartheta}$. Choosing a transformation $t(x,\boldsymbol{\vartheta})$ that creates a positive correlation between these estimators $\hat{g}_{N}(x)$ is desirable because it reduces the variance of their difference. Indeed, when comparing two first-stage decisions $x_{1}$ and $x_{2}$, the variance of the difference $\hat{g}_{N^{\prime}}(x_{1})-\hat{g}_{N^{\prime}}(x_{2})$ when both estimators use the same random samples of $\boldsymbol{\vartheta}$ is $Var\boldsymbol{(}\hat{g}_{N{{}^{\prime}}}(x_{1})\boldsymbol{)}+Var\boldsymbol{(}\hat{g}_{N{{}^{\prime}}}(x_{2})\boldsymbol{)}-2Cov\boldsymbol{(}\hat{g}_{N{{}^{\prime}}}(x_{1}),\hat{g}_{N{{}^{\prime}}}(x_{2})\boldsymbol{)}$, where $Var(\cdot)$ and $Cov(\cdot,\cdot)$ denote the variance and covariance. This is the reasoning behind the well-known efficiency-improvement technique of common random numbers (Glasserman and Yao 1992). This discussion extends to the application of any sampling technique for solving the transformed program. Section 4 further discusses this question.

### v4

@@ 2.2 Literature Review
For discrete random vectors, the elements are usually assumed to be independent. For the discrete selection of distributions, distinct applications of two-stage programs have been studied, such as network design protection (Viswanath et al. 2004, Flach 2010, Peeta et al. 2010, Prestwich et al. 2014), facility protection (Medal et al. 2016), facility location-protection (Li et al. 2021), fire risk management (Bhuiyan et al. 2019), supplier risk management (Zhou et al. 2022), facility protection and network design (Bhuiyan et al. 2020), and airline flight network expansion problems (Şafak et al. 2022). All these works used decision-dependent probabilities as in Equation (3), except for Şafak et al. (2022), who chose the probabilities of the entire random vector. These applications usually lead to large-scale instances with numerous scenarios.
@@
Prestwich et al. (2014) and Bhuiyan et al. (2019) respectively obtained exact MILP reformulations of the DE-OP by using a linear scaling procedure and the probability chain technique described by O’Hanley et al. (2013). To solve large-scale problems, Prestwich et al. (2014) proposed a bundle method, and Medal et al. (2016), Bhuiyan et al. (2020), and Zhou et al. (2022) implemented the L-shaped decomposition for solving reformulations resulting from the probability chain technique. Usually, $T_{k}$ is a zero matrix in these two-stage programs (i.e., the second stage can be solved a priori). The exception is Zhou et al. (2022), who used an exact McCormick envelope (McCormick 1976) to obtain an MILP. These reformulations apply specifically to the case of the discrete selection of distributions and rely on the binary variables in $\hat{p}_{\hat{r}(u,k)}^{u}(x)$. However, they require numerous variables and big-M constraints. Thus, the practical use of these reformulations is limited to problems with a sufficiently small number of scenarios. Flach (2010) proposed a different approximate MILP reformulation for the DE-OP by using the binary variables in $\hat{p}_{\hat{r}(u,k)}^{u}(x)$ to reformulate $p_{k}(x)$ as a set of exponential and linear constraints. The authors ensured a predefined precision by approximating the exponential constraints by piecewise linear cuts included in a branch-and-cut scheme.
@@
Various decision-dependent probability mass functions have also been studied, such as the binomial distribution (Karaesmen and Van Ryzin 2004, Green et al. 2013), linear interpolation (Du and Peeta 2014), a linear scaling and convex combination of a set of distributions (Hellemo et al. 2018), and the exponential function (Bhuiyan 2021). In most studies, the endogenous uncertainty in general cases of discrete random variables resulted in DE-OPs, which are mixed-integer nonlinear nonconvex programs (MINLPs) for which proving global optimality is harder than for MILPs or mixed-integer convex programs (MICPs). Even if $p_{k}(x)$ are convex or concave (see, e.g., Hellemo et al. 2018) or $\hat{p}_{\hat{r}(u,k)}^{u}(x)$ are convex or concave (see, e.g., Du and Peeta 2014, Bhuiyan 2021, Holzmann and Smith 2021, Homem-de Mello et al. 2022), the term $p_{k}(x)q_{k}^{\top}y_{k}$ could lead to an MINLP. Within this context, Krasko and Rebennack (2017) studied a two-stage preparedness hazard problem for debris-flow-hazard mitigation and solved the DE-OP with the BARON solver. Hellemo et al. (2018) solved the DE-OP of an energy investment problem using a generalized Benders decomposition (GBD) and the BARON solver. However, these approaches are limited to solving instances with few scenarios because the DE-OP in extensive form explicitly enumerates all scenarios.
@@
Viswanath et al. (2004), Peeta et al. (2010), and Du and Peeta (2014) used the SAA within iterative heuristic procedures to evaluate current first-stage solutions for two-stage network protection problems. Unfortunately, these procedures do not provide performance guarantees. The SAA was also used to estimate the performance of current solutions within stochastic gradient methods for overbooking (Karaesmen and Van Ryzin 2004) and appointment-scheduling problems (Homem-de Mello et al. 2022)Homem-de Mello et al. (2022) proved the local convergence of the algorithm, which is compatible with its nonconvexity. Finally, by considering a specific gradient estimator, Karaesmen and Van Ryzin (2004) proved the local convergence of the algorithm.
@@
To the best of our knowledge, only Ernst and Powell (1995), Ekin et al. (2017), Ekin (2018), and Hellemo et al. (2018) investigated continuous endogenous random variables. Ernst and Powell (1995) proposed a search procedure for solving a newsvendor problem with uncertain demand following a normal distribution with mean and variance selected by first-stage decisions. The augmented probability simulation (APS), in which decisions are also simulated, was used by Ekin et al. (2017) and Ekin (2018) for solving, respectively, a newsvendor problem with uncertain demand depending on stock level decisions and a production-planning problem with random yield depending on maintenance and production decisions. In both works, the dependency was modeled with a normal distribution with mean and variance defined according to the first-stage decisions. Finally, Hellemo et al. (2018) used the discretized function of the Kumaraswamy distribution and approximated the discretized function of the normal distribution. The discretization trades an increase in the number of scenarios for a confident discrete approximation. Approximations of the discretized function may also be needed when the cumulative distribution has no closed form or a complex closed form. In some applications, the search procedure cannot analyze the entire decision space, implying the need for a heuristic procedure. Finally, the APS remains limited to problems where the second stage is represented by analytical expressions.
@@ 3 Methodology
Our methodology is based on using random variable transformation, which is a classical topic in probability and statistics (see Hogg et al. 1977, Chap. 5), to model endogenous uncertainty. This section outlines Assumption 3, which is necessary for applying our methodology. Section 4 shows that this assumption is not restrictive by evidencing many cases in which it holds. {assumption} There exists (1) a random vector $\vartheta^{\top}(\omega)=\bm{(}\widetilde{q}(\omega)^{\top},\widetilde{h}(\omega)^{\top},\widetilde{T}(\omega),\widetilde{W}(\omega)\bm{)}$ with a known probability distribution that is independent of the first-stage decisions $x\in\mathcal{X}$ and (2) a transformation function $t(x,\bm{\vartheta})$, such that $\bm{\xi}_{x}=t\left(x,\bm{\vartheta}\right)$. Assumption 3 allows us to mimic the endogenous uncertainty by using a function $t(\cdot,\cdot)$ that combines the first-stage variables $x$ and the decision-independent random vector $\bm{\vartheta}$. For example, $t(\cdot,\cdot)$ can be the inverse transformation of $\bm{\xi}_{x}$ (Robert and Casella 1999, Chap. 2). Here, the random vector $\bm{\vartheta}$ does not necessarily have the same dimension as $\bm{\xi}$ or $\bm{\xi}_{x}$. Furthermore, the random elements of $\bm{\xi}_{x}$ may be dependent. For the sake of simplicity, we consider that $t(x,\bm{\vartheta})=(t_{1}\bm{(}x,\widetilde{q}(\omega)^{\top}\bm{)},t_{2}\bm{(}x,\widetilde{h}(\omega)\bm{)},t_{3}\bm{(}x,\widetilde{T}(\omega)\bm{)},t_{4}\bm{(}x,\widetilde{W}(\omega)\bm{)})$, where $t_{1}\bm{(}x,\widetilde{q}\left(\omega\right)^{\top}\bm{)}$, $t_{2}\bm{(}x,\widetilde{h}\left(\omega\right)\bm{)}$, $t_{3}\bm{(}x,\widetilde{T}\left(\omega\right)\bm{)}$, and $t_{4}\bm{(}x,\widetilde{W}\left(\omega\right)\bm{)}$ are transformation functions applied to the corresponding subvectors and matrices of $\bm{\vartheta}$. These functions return, respectively, the subvectors and matrices of $\bm{\xi}_{x}$, i.e., $q_{x}(\omega)^{\top}$, $h_{x}(\omega)$, $T_{x}(\omega)$, and $W_{x}(\omega)$. From Assumption 3, $t(.)$ and thus $t_{1}(.)$, $t_{2}(.)$, $t_{3}(.)$, and $t_{4}(.)$ must account for any dependency between the random elements of $\bm{\xi}_{x}$. Section 4 details different methods to define appropriate transformations respecting Assumption 3.
@@
In the field of stochastic optimization, Varaiya and Wets (1989) and Dupacová (2006) introduced the idea of transforming a decision-dependent probability distribution into an independent distribution in what is called “plug-in” or “change of measure” (see Pflug 2012 for a detailed explanation) and might employ random variable transformation. Despite these initial works, no study has directly modeled endogenous uncertainty as a function $t(.)$ nor investigated appropriate transformations for random variables following classical endogenous probability distributions. Specifically, the definition of $t(.)$ remains unexplored. Moreover, no study has focused on the nature and solution of problems resulting from this transformation nor, notably, addressed their potential nonconvexity. Finally, Holzmann and Smith (2021) used random variable transformation for a specific endogenous discrete random variable, for which Section 4.1.1 highlights the similarities and differences with the approach proposed herein. Therefore, the present work proposes a general framework using functions of random variables for addressing 2-SMIPD given Assumption 3. We also investigate practical transformations for endogenous random variables following classical probability distributions, as well as the convexity of the resulting transformed program.
Based on Assumption 3, we can formulate 2-SMIPD as the transformed program
@@
where the transformed second stage program $Q_{T}\bm{(}x,\vartheta(\omega)\bm{)}$ is defined as
$\min\{t_{1}\bm{(}x,\widetilde{q}\left(\omega\right)^{\top}\bm{)}y(\omega):t_{3}\bm{(}x,\widetilde{T}\left(\omega\right)\bm{)}x+t_{4}\bm{(}x,\widetilde{W}\left(\omega\right)\bm{)}y(\omega)\geq t_{2}\bm{(}x,\widetilde{h}\left(\omega\right)\bm{)},y(\omega)\in\mathcal{Y}\},$ (5)
@@
Under Assumption 3, $g(x)=\mathbb{E}_{\bm{\xi}\mid x}\left[Q\bm{(}x,\xi(\omega)\bm{)}\right]=\mathbb{E}_{\bm{\vartheta}}\left[Q_{T}\bm{(}x,\vartheta(\omega)\bm{)}\right]$ and the 2-SMIPD (Program (1)) can be reformulated into the transformed Program (4). In other words, Programs (1) and (4) have the same optimal solutions.
Thus, our methodology is established on finding an appropriate transformation function to obtain Program (5). In that way, under Assumption 3, 2-SMIPD can be transformed into a two-stage stochastic program with exogenous uncertainty. Solution procedures for the transformed Program (4) based on its properties are well-studied (see, e.g., Shapiro et al. 2021) and, thus, can be applied in this case. Program (5) has the terms $t_{1}\bm{(}x,\widetilde{q}\left(\omega\right)^{\top}\bm{)}y$, $t_{2}\bm{(}x,\widetilde{h}\left(\omega\right)\bm{)}$, $t_{3}\bm{(}x,\widetilde{T}\left(\omega\right)\bm{)}x$, and $t_{4}\bm{(}x,\widetilde{W}\left(\omega\right)\bm{)}y$, which might be nonlinear nonconvex. We then introduce the following assumption: {assumption} The continuous relaxation of the second-stage Program (5) is convex. Program (5) might need to be reformulated to convexify its continuous relaxation. Section 4 presents different cases for which Assumption 3 holds. In addition, the exogenous random vector $\bm{\vartheta}$ might have a large or infinite dimension (see, e.g., the inversion in Section 4.1). Hence, next, we describe the application of the SAA (Shapiro et al. 2021, Chap. 5) to solve Program (4), which, although intuitive, shows that the transformation does not compromise desirable SAA properties. Nonetheless, other procedures can also be employed to solve Program (4) in this case, e.g., stochastic gradient (Shapiro et al. 2021).
@@ 3.1 Sample Average Approximation
We consider $N$ independent and identically distributed samples $(\bm{\vartheta^{1}},\ldots,\bm{\vartheta^{N}})$ of $\bm{\vartheta}$ with realizations $(\vartheta^{1},\ldots,\vartheta^{N})$. The transformed SAA program is
@@
where $\hat{g}_{N}(x)$ is the SAA estimator of $\mathbb{E}_{\bm{\vartheta}}\left[Q_{T}\bm{(}x,\vartheta(\omega)\bm{)}\right]$ defined as
@@
This estimator converges point-wise with probability one (w.p.1) to $\mathbb{E}_{\bm{\vartheta}}[Q_{T}\bm{(}x,\vartheta(\omega)\bm{)}]$ when ${N\rightarrow+\infty}$. From Lemma 3.1, we have $\mathbb{E}[\hat{g}_{N}(x)]=g(x)$, so $\hat{g}_{N}(x)$ is an unbiased estimator of $g(x)$. We denote $v^{*}$, $v_{t(x,\bm{\vartheta})}^{*}$ and $v_{N}$ as the optimal values of Program (1), Program (4), and Program (6), respectively. Under some mild conditions, such as when Program (5) is convex, an MILP or when the first-stage set $\mathcal{X}$ is finite, $v_{N}$ converges to $v_{t(x,\bm{\vartheta})}^{*}$ w.p.1 as ${N\rightarrow+\infty}$ (cf. Shapiro et al. 2021, Chap. 5, and Kleywegt et al. 2002, respectively). In these cases and under Assumption 3, $v_{N}$ converges to $v^{*}$ w.p.1 as $N\rightarrow+\infty$, since $v^{*}=v_{t(x,\bm{\vartheta})}^{*}$ from Lemma 3.1. This result, and its rate of convergence (cf. (Shapiro et al. 2021)), depends on the properties of $\bm{\vartheta}$ (i.e., on the convergence of $v_{N}$ to $v_{t(x,\bm{\vartheta)}}$), which may have larger dimension and support than $\bm{\xi}$ and may require a larger number of samples.
@@
The SAA finds a feasible first-stage solution $\overline{x}\in\mathcal{X}$. For that, we solve $M$ instances of the SAA Program (6), each with $N$ independent samples. We then evaluate the optimal solutions of every instance by using a larger sample size $N^{\prime}$ and define $\overline{x}$ as the optimal solution with the best evaluation. When $Q\bm{(}x,\xi(\omega)\bm{)}$ is finite and Program (1) has relatively complete recourse (i.e., the second stage associated with $x\in\mathcal{X}$ is feasible for almost every realization of $\bm{\xi}$), then $Q_{T}\bm{(}x,\vartheta(\omega)\bm{)}$ is also finite, and Program (4) also has relatively complete recourse, given that $\bm{\xi}_{x}=t(x,\bm{\vartheta})$ from Assumption 3. In this case, in addition to finding $\overline{x}$, we can compute a lower-bound estimator $\overline{v}_{M}^{N}$ of $v^{*}$, an upper-bound estimator $\hat{v}_{N^{\prime}}(\overline{x})$ of $v^{*}$, an over-estimator $\overline{gap}(\overline{x})$ of the optimality gap $c^{\top}\overline{x}+g(\overline{x})-v^{*}$, and unbiased estimators of the variances of $\overline{v}_{N}^{M}$, $\hat{v}_{N{{}^{\prime}}}(\overline{x})$, and $\overline{gap}(\overline{x})$, i.e., $\hat{\sigma}_{N,M}^{2}$, $\hat{\sigma}_{N{{}^{\prime}}}^{2}(\overline{x})$, and $\hat{\sigma}_{gap}^{2}(\overline{x})$, respectively. The Appendix 7 details these estimators and the SAA’s procedure.
@@
By defining the $\bm{\xi}_{x}$ with the function $t(x,\bm{\vartheta})$, the estimators $\hat{g}_{N}(x)$ for all $x\in\mathcal{X}$ are based on the same samples of $\bm{\vartheta}$. Choosing a transformation $t(x,\bm{\vartheta})$ that creates a positive correlation between these estimators $\hat{g}_{N}(x)$ is desirable because it reduces the variance of their difference. Indeed, when comparing two first-stage decisions $x_{1}$ and $x_{2}$, the variance of the difference $\hat{g}_{N}(x_{1})-\hat{g}_{N}(x_{2})$ when both estimators use the same random samples of $\bm{\vartheta}$ is $Var\bm{(}\hat{g}_{N}(x_{1})\bm{)}+Var\bm{(}\hat{g}_{N}(x_{2})\bm{)}-2Cov\bm{(}\hat{g}_{N}(x_{1}),\hat{g}_{N}(x_{2})\bm{)}$, where $Var(\cdot)$ and $Cov(\cdot,\cdot)$ denote the variance and covariance. This is the reasoning behind the well-known efficiency-improvement technique of common random numbers (Glasserman and Yao 1992). This discussion extends to the application of any sampling technique for solving the transformed program. Section 4 further discusses this question.